OpenStudy (xbeenchantedx):

A runaway train takes 12 seconds to stop after the emergency brakes are applied. If the emergency brakes can slow the train at the rate of 0.5 m/s2, how fast was the train going before the brakes were applied?

OpenStudy (anonymous):

t=12s v=0 m/s (Stops) a=-5 m/s2 (Deceleration) u=? v=u+at 0=u-(5*12) u=5*12 u=60 m/s